# Cool-Core Destruction in Merging Clusters with AGN Feedback and Radiative Cooling

### v1

@@ 5 Conclusions
For our $mcf$ simulations (i.e., mergers with AGN feedback and radiative cooling), we examine the entropy map and the evolution of central entropy, as well as heating and cooling powers and energies. In the mass ratio $R=1$ cases, gas mixing is more thorough in off-axis mergers, leading to a larger entropy increase. Except for the head-on merger, all final remnants are NCC clusters. Conversely, in the $R=3$ cases, head-on mergers more effectively disrupt the primary core, resulting in a larger entropy increase. Nonetheless, all clusters eventually transform into NCCs. Notably, in these off-axis mergers, the jet power is strong in the early stages due to late mergers. The gas inflow caused by sloshing maintains this strong jet power, leading to overheating of the cluster cores. In the $R=10$ cases, the disruption from the secondary cluster is insufficient to destroy the CCs, and the balance between heating and cooling resembles the single cluster case.
@@
Special thanks to Chun-Yen Chen for his assistance in reviewing the simulation source code and resolving bugs, which improved the code’s performance. SSC acknowledges support from the Research Grant for University Students (111-2813-C-002-008-M and 112-2813-C-002-072-M), sponsored by the National Science and Technology Council in Taiwan. HYKY acknowledges support from National Science and Technology Council (NSTC) of Taiwan (NSTC 112-2628-M-007-003-MY3) and Yushan Scholar Program of the Ministry of Education (MoE) of Taiwan (MOE-108-YSFMS-0002-003-P1). HYS acknowledges funding support from NSTC of Taiwan under Grants No. NSTC 111-2628-M-002-005-MY4 and the NTU Academic Research-Career Development Project under Grant No. NTU-CDP-113L7729. JAZ is funded by the Chandra X-ray Center, which is operated by the Smithsonian Astrophysical Observatory for and on behalf of NASA under contract NAS8-03060. MG acknowledges funding support from the ERC Consolidator Grant BlackHoleWeather (101086804). This work uses high-performance computing facilities operated by the Computational Astrophysics Lab (CALab) at NTU. Data analysis presented in this paper is conducted with the publicly available yt visualization software (Turk et al., 2010).
@@ Appendix A High-Resolution Convergence Test
In order to verify the convergence of our simulations, we performed a high-resolution run of the $mcf5$ simulation. In this run, the resolution for the central 20 kpc was doubled compared to the standard run, resulting in a spatial resolution of $\Delta x=$ 0.46 kpc. Additionally, the particle number was increased by a factor of five. At the end of the simulation, the radial entropy profiles of both the standard and high-resolution runs are in excellent agreement. Figure 9 illustrates the comparison between the high-resolution run and the standard run, with the left panel showing the central entropy $S_{40}$ and the right panel depicting the heating/cooling rates. Overall, the evolution of the standard and high-resolution runs is consistent. While transient feedback and cooling cycles exhibit sensitivity to numerical errors, causing discrepancies during 3-4 Gyr, these differences do not impact the final states of the clusters. This confirms that our simulations are numerically converged.

### v3

@@ V Conclusions
For our mcfmcf simulations (i.e., mergers with AGN feedback and radiative cooling), we examine the entropy map and the evolution of central entropy, as well as heating and cooling powers and energies. In the mass ratio $R=1$ cases, gas mixing is more thorough in off-axis mergers, leading to a larger entropy increase. Except for the head-on merger, all final remnants are NCC clusters. In the $R=3$ cases, the outcomes are more sensitive to the intricate interplay among mergers, cooling and AGN feedback. In particular, head-on mergers ($mcf4$) can effectively disrupt the primary core, resulting in a NCC cluster. For off-axis mergers, the CC could either be preserved throughout the simulation ($mcf5$) or be disrupted ($mcf6$), depending on whether the AGN could effectively preheat the core prior to the infall of the seconary cluster. In the $R=10$ cases, the disruption from the secondary cluster is insufficient to destroy the CCs, and the balance between heating and cooling resembles the single cluster case.
@@
SSC acknowledges support from the Research Grant for University Students (111-2813-C-002-008-M and 112-2813-C-002-072-M), sponsored by the National Science and Technology Council in Taiwan. HYKY acknowledges support from National Science and Technology Council (NSTC) of Taiwan (NSTC 112-2628-M-007-003-MY3; NSTC 114-2112-M-007-032-MY3) and Yushan Scholar Program of the Ministry of Education (MoE) of Taiwan (MOE-108-YSFMS-0002-003-P1; MOE-114-YSFMS-0002-002-P2). HYS acknowledges funding support from NSTC of Taiwan under Grants No. NSTC 111-2628-M-002-005-MY4 and the NTU Academic Research-Career Development Project under Grant No. NTU-CDP-113L7729. JAZ is funded by the Chandra X-ray Center, which is operated by the Smithsonian Astrophysical Observatory for and on behalf of NASA under contract NAS8-03060. MG acknowledges funding support from the ERC Consolidator Grant BlackHoleWeather (101086804). This work uses high-performance computing facilities operated by the Computational Astrophysics Lab (CALab) at NTU. Data analysis presented in this paper is conducted with the publicly available yt visualization software (Turk et al., 2010).
@@ Appendix A High-Resolution Convergence Test
In order to verify the convergence of our simulations, we performed a high-resolution run of the $mcf6$ simulation. In this run, the resolution for the central 20 kpc was doubled compared to the standard run, resulting in a spatial resolution of $\Delta x=$ 0.46 kpc. Additionally, the particle number was increased by a factor of five. At the end of the simulation, the radial entropy profiles of both the standard and high-resolution runs are in excellent agreement. Figure 9 illustrates the comparison between the high-resolution run and the standard run, with the left panel showing the central entropy $S_{40}$ and the right panel depicting the heating/cooling rates. Overall, the evolution of the standard and high-resolution runs is consistent. In the high-resolution run, the AGN feedback is more prominent due to formation of more dense, cold clumps. The more powerful jet activities during t=34t=3-4 Gyr more efficiently preheat the CC and elevate the core entropy, reaching a higher entropy level compared to the simulation with fiducial resolution. Despite this difference, the simulation outcome that the CC is destroyed by the combined effect of merger and AGN is robust.